# Constraining the gravitational-wave emission of core-collapse supernovae with ground-based detectors

### v1

@@ IV Detection prospects
Figure 4: The signal to noise ratio of the GWB $(S/N)_{\rm GWB}$ (solid lines with shaded band) or the detection rate $R_{\rm det}$ (dashed lines with shaded band) of individual CCSNe events as a function the GW energy emitted by CCSNe events for the ET (left panel) or CE (right panel) detector. Blue and red curves denote two CCSNe spectral models (see text for details): Ott+11 s27-fheat1.05 (blue) and And+19 m15r (red). The crossing between the $(S/N)_{\rm GWB}$ lines and the $R_{\rm det}$ lines with their respective thresholds (horizontal lines) are marked with circles and squares (with the GW energy values listed), respectively. Two vertical lines are the upper limits derived in this work.
In this section, we explore the detection prospects of both the GWB and individual CCSNe events for future detectors. Assuming Gaussian detector noise, the optimal signal-to-noise ratio ($S/N$) for a cross-correlation analysis over an observation time $T$ is given by Allen and Romano (1999):
@@
where $P_{1}(f)$ and $P_{2}(f)$ denote the power spectral density (PSD) of the two detectors, and $\gamma(f)$ is the overlap reduction function (ORF). In this work, we adopt a detection threshold of $(S/N)_{\mathrm{GWB}}=3$, accumulated over one year of observation, to assess the detectability of the GWB from CCSNe.
We further evaluate the detectability of individual CCSNe events using the matched-filter method, assuming optimal detector alignment and isotropic GW emission. The matched-filter $S/N$ can be expressed as Flanagan and Hughes (1998); Andresen et al. (2019):
@@
$h_{c}(f)=\sqrt{\frac{2G}{\pi^{2}c^{3}D_{L}^{2}(z)}\frac{\mathrm{d}E}{\mathrm{d}f}}.$ (10)
@@
To calculate the detectable event rate ($R_{\text{det}}$), we follow the approach in Ref. Szczepanczyk et al. (2021), adopting a matched-filter detection threshold of $(S/N)_{\mathrm{CCSNe}}=20$ and assuming a detection efficiency of $50\%$. We consider three detector configurations, A+ LIGO Scientific Collaboration (2018), the Einstein Telescope (ET) Hild et al. (2011) and the Cosmic Explorer (CE) Evans et al. (2021). Specifically, we consider the ET-D configuration with three nearly co-located detectors, and a pair of two detectors with $40~\rm{km}$ and $20~\rm{km}$ arm lengths for CE. For simplicity, we refer to these respective detector pairs as ET and CE throughout this work. The sensitivity curves for these detectors are obtained from Ref. Abbott et al. (2017). Details of the detector configurations and the corresponding ORFs can be found in Ref. Gupta et al. (2024). The ORFs used in our analysis are computed following Ref. Allen et al. (2002).
@@
Figure 3 illustrates the detection prospects for the GWB (top panels) or individual CCSNe events (bottom panels) for three future detectors: A+ (left), CE (middle), and ET (right), assuming Gaussian emission spectrum as described by Eq. (3). Each panel is color-coded by $\log_{10}(E_{\mathrm{GW}}/M_{\odot}c^{2})$, the required GW energy emitted a single CCSNe event, as a function of the peak frequency $f_{\mathrm{peak}}$ and the spectral width $\Delta$, to achieve a GWB detection with $(S/N)_{\mathrm{GWB}}=3$ or a detection rate of one CCSNe event per year ($R_{\mathrm{det}}=1$).
@@
The GW energy required for detection spans a wide range—typically over 2 to 3 orders of magnitude—depending on the choice of spectral parameters and detector sensitivity. Notably, the detectability is strongly dependent on both $f_{\mathrm{peak}}$ and $\Delta$. We find that high-frequency, narrow-band signals (e.g., $f_{\mathrm{peak}}\gtrsim 1000~\mathrm{Hz}$, $\Delta\lesssim 500~\mathrm{Hz}$) require significantly more GW energy to be detectable. This is due to the increasing detector noise at high frequencies, and in the case of GWB detection the reduced cross-correlation sensitivity defined by the overlap reduction function. In comparison, low-frequency signals ($f_{\mathrm{peak}}\lesssim 200~\mathrm{Hz}$) allow for detection with GW energy $E_{\mathrm{GW}}$ lower by a factor of 10 to 100. Across three detectors, we can see that the required GW energy for the detection of the GWB is generally higher than that for a detection rate of one single CCSNe event per year. This indicates that individual CCSNe events are likely to be detected prior to their cumulative GWB becoming observable.
@@
Each panel of Figure 3 includes several red dashed contours, indicating constant levels of $E_{\mathrm{GW}}$. For the A+ panels, we include two lines corresponding to upper limits derived in this work for Prior I ($E_{\mathrm{GW}}=9.9\times 10^{-3}~M_{\odot}c^{2}$) and Prior II ($E_{\mathrm{GW}}=3.2\times 10^{-2}~M_{\odot}c^{2}$). As a guideline, we can see that the A+ detector would be able to further improve upper limits on $E_{\mathrm{GW}}$ to $10^{-3}$ and $10^{-4}~M_{\odot}c^{2}$ levels for low-frequency signals ($f_{\mathrm{peak}}\lesssim 200~\mathrm{Hz}$), through the GWB and single-event searches, respectively. For third-generation detectors such as ET and CE, the detection prospects are more promising. There are chances for detection of the GWB if the GW energy is above $10^{-4}~M_{\odot}c^{2}$ for a large fraction of the parameter space. For single events, the detection limit can be pushed to $10^{-5}~M_{\odot}c^{2}$ or even $10^{-6}~M_{\odot}c^{2}$ for low-frequency signals.
@@
So far our calculations are based on the assumption that the GW emission spectrum from CCSNe events can be approximated by a Gaussian function. To account for uncertainties in the spectral shape of GW emission from CCSNe events, we consider two CCSNe spectral models from Fig. 3 of Ref. Szczepanczyk et al. (2021): the Ott+11 s27-fheat1.05 model (blue) and the And+19 m15r model (red). We keep the spectral shape of these two models and rescale the amplitude [^1] This is for illustrative purpose only as we aim to find a plausible range of required GW energy for detection of GWs from CCSNe with third-generation detectors., to calculate the GWB single-to-noise ratio $(S/N)_{\mathrm{GWB}}$ and single-event detection rate $R_{\text{det}}$ for ET and CE. Results are shown in Fig. 4. Solid lines indicate the $(S/N)_{\mathrm{GWB}}$ as a function of the average GW energy per CCSNe, while dashed lines show the corresponding detectable event rate per year. The purple and orange shaded bands represent model-dependent uncertainties. In the same figure, the gray and gold dashed horizontal lines mark the detection thresholds $(S/N)_{\mathrm{GWB}}=3$ and $R_{\text{det}}=1$, respectively. The vertical lines correspond to the GW energy constraints obtained in this work.
@@
In Fig. 4, one can draw consistent conclusions as in Figure 3. First, the requirement on the GW energy for a detection of the GWB is generally higher than that for individual CCSNe events. This is most significant for CE, with two differing by a factor of 10, whereas the difference is around $3$ for ET. This can be attributed to differences in detector noise characteristics and the ORF. The GWB can be detectable with third-generation detectors if the GW energy is $\gtrsim 10^{-4}~M_{\odot}c^{2}$, whereas a detection of single CCSNe event is possible if $E_{\text{GW}}\gtrsim 10^{-5}~M_{\odot}c^{2}$.
@@ V Conclusions
Core-collapse supernovae (CCSNe) have long been considered to be an important source of GWs. While direct detections from CCSNe remain elusive, constraining their average GW energy emission is crucial for evaluating the detectability of CCSNe signals. In this work, we use cross-correlation measurements from the third observing run (O3) of LIGO and Virgo detectors to place upper limits on the average GW energy emitted per CCSNe event [^2] At the time of writing this paper, results of a GWB search using the first part of O4 data were released Abac et al. (2025). We expect that our upper limits would be slightly improved if O4a data were used.. By adopting a generic Gaussian spectrum model, we drive constraints that are marginalized over a wide range of emission peak frequency and spectral width. Furthermore, we assess the detection prospects with third-generation ground-based detectors,including ET and CE, for both individual events and the GWB produced by a cosmic population of CCSNe sources.
@@
Finally, we compare the detectability of the GWB and individual CCSNe events for future detectors. In general, individual events are expected to be detectable prior to the GWB. For third-generation detectors such as ET and CE, we find that the GWB can be detectable if the GW energy is $\gtrsim 10^{-4}~M_{\odot}c^{2}$, and a detection of a single CCSNe event is possible within one year of operation if $E_{\text{GW}}\gtrsim 10^{-5}~M_{\odot}c^{2}$. We also find that for GW emission with a peak frequency above $1000\,\mathrm{Hz}$, the sensitivity of ET or CE to both the GWB and single events decreases by a factor of $\sim 10$ in comparison to signals with peak frequencies below $500\,\mathrm{Hz}$. This highlights the scientific potential of dedicated kHz detectors Miao et al. (2018); Martynov et al. (2019); Ackley et al. (2020); Zhang et al. (2023).
@@ Acknowledgements.
This work is supported by the National Key Research and Development Program of China (No. 2023YFC2206704), the Fundamental Research Funds for the Central Universities, and the Supplemental Funds for Major Scientific Research Projects of Beijing Normal University (Zhuhai) under Project ZHPT2025001. This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan.

### v2

@@ IV Detection prospects
Figure 4: The signal to noise ratio of the GWB $(S/N)_{\rm GWB}$ (solid lines with shaded band) or the detection rate $R_{\rm det}$ (dashed lines with shaded band) of individual CCSNe events as a function the GW energy emitted by CCSNe events for the ET (left) or CE (right) detector. Blue and red curves denote two CCSNe spectral models (see text for details): Ott+11 s27-fheat1.05 (blue) and And+19 m15r (red). The crossing between the $(S/N)_{\rm GWB}$ lines and the $R_{\rm det}$ lines with their respective thresholds (horizontal lines) are marked with circles and squares (with the GW energy values listed), respectively. Two vertical lines are the upper limits derived in this work.
In this section, we explore the detection prospects of both the GWB and individual CCSNe events for future detectors. Assuming Gaussian detector noise, the optimal signal-to-noise ratio ($S/N$) for a cross-correlation analysis over an observation time $T$ is given by Allen and Romano (1999)
@@
where $P_{1}(f)$ and $P_{2}(f)$ denote the power spectral density of the two detectors, and $\gamma(f)$ is the overlap reduction function (ORF). In this work, we adopt a detection threshold of $(S/N)_{\mathrm{GWB}}=3$, accumulated over one year of observation, to assess the detectability of the GWB from CCSNe.
We further evaluate the detectability of individual CCSNe events using the matched-filter method, assuming optimal detector alignment and isotropic GW emission. The matched-filter $S/N$ can be expressed as Flanagan and Hughes (1998); Andresen et al. (2019)
@@
$h_{c}(f)=\sqrt{\frac{2G}{\pi^{2}c^{3}D_{L}^{2}(z)}\frac{\mathrm{d}E}{\mathrm{d}f_{s}}}.$ (10)
@@
To calculate the detectable event rate ($R_{\text{det}}$), we follow the approach in Ref. Szczepanczyk and others (2021), adopting a matched-filter detection threshold of $(S/N)_{\mathrm{CCSNe}}=20$ and assuming a detection efficiency of $50\%$. We consider three detector configurations: A+ LIGO Scientific Collaboration (2018), the Einstein Telescope (ET) Hild and others (2011), and the Cosmic Explorer (CE) Evans and others (2021). Specifically, we consider the ET-D configuration with three nearly colocated detectors, and a pair of two detectors with $40$ and $20~\rm{km}$ arm lengths for CE. For simplicity, we refer to these respective detector pairs as ET and CE throughout this work. The sensitivity curves for these detectors are obtained from Ref. Abbott and others (2017). Details of the detector configurations and the corresponding ORFs can be found in Ref. Gupta and others (2024). The ORFs used in our analysis are computed following Ref. Allen et al. (2002).
@@
Figure 3 illustrates the detection prospects for the GWB (top panels) or individual CCSNe events (bottom panels) for three future detectors: A+ (left), CE (middle), and ET (right), assuming Gaussian emission spectrum as described by Eq. (3). Each panel is color coded by $\log_{10}(E_{\mathrm{GW}}/M_{\odot}c^{2})$, the required GW energy emitted a single CCSNe event, as a function of the peak frequency $f_{\mathrm{peak}}$ and the spectral width $\Delta$, to achieve a GWB detection with $(S/N)_{\mathrm{GWB}}=3$ or a detection rate of one CCSNe event per year ($R_{\mathrm{det}}=1$).
@@
The GW energy required for detection spans a wide range—typically over 2-3 orders of magnitude—depending on the choice of spectral parameters and detector sensitivity. Notably, the detectability is strongly dependent on both $f_{\mathrm{peak}}$ and $\Delta$. We find that high-frequency, narrow-band signals (e.g., $f_{\mathrm{peak}}\gtrsim 1000$, $\Delta\lesssim 500~\mathrm{Hz}$) require significantly more GW energy to be detectable. This is due to the increasing detector noise at high frequencies and, in the case of GWB detection, the reduced cross-correlation sensitivity defined by the overlap reduction function. In comparison, low-frequency signals ($f_{\mathrm{peak}}\lesssim 200~\mathrm{Hz}$) allow for detection with GW energy $E_{\mathrm{GW}}$ lower by a factor of 10-100. Across three detectors, we can see that the required GW energy for the detection of the GWB is generally higher than that for a detection rate of one single CCSNe event per year. This indicates that individual CCSNe events are likely to be detected prior to their cumulative GWB becoming observable.
@@
Each panel of Fig. 3 includes several red dashed contours, indicating constant levels of $E_{\mathrm{GW}}$. For the A+ panels, we include two lines corresponding to upper limits derived in this work for Prior I ($E_{\mathrm{GW}}=9.9\times 10^{-3}~M_{\odot}c^{2}$) and Prior II ($E_{\mathrm{GW}}=3.2\times 10^{-2}~M_{\odot}c^{2}$). As a guideline, we can see that the A+ detector would be able to further improve upper limits on $E_{\mathrm{GW}}$ to $10^{-3}$ and $10^{-4}~M_{\odot}c^{2}$ levels for low-frequency signals ($f_{\mathrm{peak}}\lesssim 200~\mathrm{Hz}$), through the GWB and single-event searches, respectively. For third-generation detectors such as ET and CE, the detection prospects are more promising. There are chances for detection of the GWB if the GW energy is above $10^{-4}~M_{\odot}c^{2}$ for a large fraction of the parameter space. For single events, the detection limit can be pushed to $10^{-5}~M_{\odot}c^{2}$ or even $10^{-6}~M_{\odot}c^{2}$ for low-frequency signals.
@@
So far our calculations are based on the assumption that the GW emission spectrum from CCSNe events can be approximated by a Gaussian function. To illustrate the impact of CCSN waveform modeling uncertainties on our detectability forecasts, we consider two CCSNe spectral models from Fig. 3 of Ref. Szczepanczyk and others (2021): the Ott+11 s27-fheat1.05 model (blue) and the And+19 m15r model (red). [^2] A comprehensive survey of all available CCSN numerical-simulation waveforms is impractical. We therefore evaluated six waveforms shown in Fig. 3 of Ref. Szczepanczyk and others (2021) and selected these two representative models with intermediate (i.e., nonextreme) results for demonstration purposes. We keep the spectral shape of these two models and rescale the amplitude [^3] This is for illustrative purpose only as we aim to find a plausible range of required GW energy for detection of GWs from CCSNe with third-generation detectors. to calculate the GWB single-to-noise ratio $(S/N)_{\mathrm{GWB}}$ and single-event detection rate $R_{\text{det}}$ for ET and CE. Results are shown in Fig. 4. Solid lines indicate the $(S/N)_{\mathrm{GWB}}$ as a function of the average GW energy per CCSNe, while dashed lines show the corresponding detectable event rate per year. The purple and orange shaded bands represent model-dependent uncertainties. In the same figure, the gray and gold dashed horizontal lines mark the detection thresholds $(S/N)_{\mathrm{GWB}}=3$ and $R_{\text{det}}=1$, respectively. The vertical lines correspond to the GW energy constraints obtained in this work.
@@
In Fig. 4, one can draw consistent conclusions as in Fig. 3. First, the requirement on the GW energy for a detection of the GWB is generally higher than that for individual CCSNe events. This is most significant for CE, with two differing by a factor of 10, whereas the difference is around $3$ for ET. This can be attributed to differences in detector noise characteristics and the ORF. The GWB can be detectable with third-generation detectors if the GW energy is $\gtrsim 10^{-4}~M_{\odot}c^{2}$, whereas a detection of single CCSNe event is possible if $E_{\text{GW}}\gtrsim 10^{-5}~M_{\odot}c^{2}$.
@@ V Conclusions
Core-collapse supernovae have long been considered to be an important source of GWs. While direct detections from CCSNe remain elusive, constraining their average GW energy emission is crucial for evaluating the detectability of CCSNe signals. In this work, we use cross-correlation measurements from the third observing run of LIGO and Virgo detectors to place upper limits on the average GW energy emitted per CCSNe event [^4] At the time of writing this paper, results of a GWB search using the first part of O4 data were released Abac and others (2025b). We expect that our upper limits would be slightly improved if O4a data were used.. By adopting a generic Gaussian spectrum model, we drive constraints that are marginalized over a wide range of emission peak frequency and spectral width. Furthermore, we assess the detection prospects with third-generation ground-based detectors,including ET and CE, for both individual events and the GWB produced by a cosmic population of CCSNe sources.
@@
Finally, we compare the detectability of the GWB and individual CCSNe events for future detectors. In general, individual events are expected to be detectable prior to the GWB. For third-generation detectors such as ET and CE, we find that the GWB can be detectable if the GW energy is $\gtrsim 10^{-4}~M_{\odot}c^{2}$, and a detection of a single CCSNe event is possible within one year of operation if $E_{\text{GW}}\gtrsim 10^{-5}~M_{\odot}c^{2}$. We also find that for GW emission with a peak frequency above $1000\,\mathrm{Hz}$, the sensitivity of ET or CE to both the GWB and single events decreases by a factor of $\sim 10$ in comparison to signals with peak frequencies below $500\,\mathrm{Hz}$. This highlights the scientific potential of dedicated kilohertz detectors Miao et al. (2018); Martynov et al. (2019); Ackley and others (2020); Zhang et al. (2023).
@@ Acknowledgements.
We thank the anonymous referees for useful comments on the manuscript. This work is supported by the National Key Research and Development Program of China (No. 2023YFC2206704), the Fundamental Research Funds for the Central Universities, and the Supplemental Funds for Major Scientific Research Projects of Beijing Normal University (Zhuhai) under Project ZHPT2025001. This research has made use of data or software obtained from the Gravitational Wave Open Science Centergwosc, a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan.
VI Data availability
The data that support the findings of this article are openly availableet al. (LIGO-Virgo-KAGRA Collaborations)
Appendix A Reproduction of LIGO-VIRGO-KAGRA (LVK) results
To validate our Bayesian inference method, we reproduce the LVK O3 isotropic stochastic background analysis presented in Ref. Abbott and others (2021b). Specifically, we analyze the same publicly released O3 cross-correlation data products and adopt the same two-parameter power-law model, $\Omega_{\rm GW}(f)=\Omega_{\rm ref}(f/f_{\rm ref})^{\alpha}$, with the same reference frequency and prior choices as in Ref. Abbott and others (2021b).
Figure 5 compares the resulting posterior distributions with those reported in Ref. Abbott and others (2021b). We find excellent agreement in both the one-dimensional marginals and the two-dimensional credible regions, with any small differences consistent with finite-sampling uncertainty (and, potentially, minor implementation details). This reproduction test provides an independent validation of our analysis framework before applying it to the CCSNe and CCSNe+CBC models considered in this work.
Figure 5: Posterior distributions of the power-law GWB model parameters $(\log_{10}\Omega_{\rm ref},\,\alpha)$ using the O3 cross-correlation data. The blue contours correspond to the results reported in Ref. Abbott and others (2021b), while the red contours show the posteriors obtained in this work using the same data products. The diagonal panels show the marginalized one-dimensional posteriors, and the lower-left panel shows the joint posterior. Solid (dashed) contours enclose the 68% (95%) credible regions.